# October 17, 2025

### v2

@@ V.9.1 Scattering Matrices
$A$ $=$ $\begin{bmatrix}2\lambda_{11,33}+2\lambda_{13,31}&2\sqrt{2}\lambda^{*}_{12,32}\\[2.84526pt] 2\sqrt{2}\lambda_{12,32}&2\lambda_{22,22}\end{bmatrix}$, (128)
$B$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda^{*}_{12,13}\\[2.84526pt] 2\sqrt{2}\lambda_{12,13}&2\lambda_{22,33}+2\lambda_{23,32}\end{bmatrix}$, (129)
$C$ $=$ $\begin{bmatrix}2\lambda_{33,33}&2\sqrt{2}\lambda_{13,23}\\[2.84526pt] 2\sqrt{2}\lambda^{*}_{13,23}&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (130)
@@
$D$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{13,23}&2\lambda^{*}_{12,32}\\ 2\lambda^{*}_{13,23}&2\lambda_{11,33}&2\lambda^{*}_{12,13}\\ 2\lambda_{12,32}&2\lambda_{12,13}&2\lambda_{11,22}\end{bmatrix}$, (132)
@@
$F$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{12,32}&2\lambda^{*}_{13,23}\\ 2\lambda^{*}_{12,32}&2\lambda_{11,22}&2\lambda^{*}_{12,13}\\ 2\lambda_{13,23}&2\lambda_{12,13}&2\lambda_{11,33}\end{bmatrix}$, (134)
@@
$G$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{13,23}&6\lambda_{12,32}\\ 6\lambda_{13,23}&2\lambda_{11,33}+4\lambda_{13,31}&6\lambda_{12,13}\\ 6\lambda^{*}_{12,32}&6\lambda^{*}_{12,13}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (136)
$H$ $=$ $\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&4\lambda_{11,33}+2\lambda_{13,31}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&4\lambda_{22,33}+2\lambda_{23,32}\\ 4\lambda_{11,33}+2\lambda_{13,31}&4\lambda_{22,33}+2\lambda_{23,32}&6\lambda_{33,33}\end{bmatrix}$, (137)
$I$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{12,32}&6\lambda_{13,23}\\ 6\lambda_{12,32}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,13}\\ 6\lambda^{*}_{13,23}&6\lambda^{*}_{12,13}&2\lambda_{11,33}+4\lambda_{13,31}\end{bmatrix}$, (138)
@@
V.9.2 Unitarity Bounds
We show the results of partial wave unitarity bounds to the eigenvalues that can be computed analytically,
@@ V.9.2 Unitarity Bounds
$\left|\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\pm\sqrt{\left(\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\right)^{2}-4\left(\lambda_{11,11}\lambda_{22,33}+\lambda_{11,11}\lambda_{23,32}-2|\lambda_{12,13}|^{2}\right)}\right|$ $\leq 8\pi,$ (143)
$\left|\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\pm\sqrt{\left(\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\right)^{2}-4\left(\lambda_{11,33}\lambda_{22,22}+\lambda_{13,31}\lambda_{22,22}-2|\lambda_{12,32}|^{2}\right)}\right|$ $\leq 8\pi,$ (144)
$\left|\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\pm\sqrt{\left(\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\right)^{2}-4\left(\lambda_{11,22}\lambda_{33,33}+\lambda_{12,21}\lambda_{33,33}-2\left|\lambda_{13,23}\right|^{2}\right)}\right|$ $\leq 8\pi.$ (145)
Term Our Notation Notation of Bento:2017eti
@@
Table 8: Comparison of coupling notation.
We compared our results with those presented in Bento:2017eti , and found them to be consistent. Table 8 provides a summary of the correspondence between the quartic couplings used in our study and those in Bento:2017eti .
@@ VI Overview
Almost all models addressing the outstanding issues in the SM include extra $SU(2)$ singlet and/or doublet scalars. In particular, many models addressing the dark matter problem include extra neutral singlet scalars. Such models must be subject to theoretical constraints, even before a simulation starts. Those constraints include boundedness from below, nonexistence of lower lying alternative vaccua, and the unitarity bounds on $2\rightarrow 2$ scattering. In this paper we address fully the perturbative partial-wave unitarity for the tree-level scattering matrix in models with any number of scalar doublets, neutral singlets, and/or charged singlets. Enforcing the correct high-energy behavior, provides bounds on the quartic couplings, freeing us from defining the exact nature of the quadratic and cubic couplings.
In contrast, if one one wishes to turn such bounds on restrictions on masses, mixing angles and other directly observable quantities, then one must define the full theory. That is, one must define the quadratic and cubic couplings, a specific vacuum, and the mass matrices must be dully diagonalized. After this, one would strife to invert the relations, turning the unitarity bounds on the quartic couplings into restrictions on combinations of masses and mixing angles. This is possible in simple models, but increasingly more difficult as the number of fields increases. In contrast, although the matrices get larger, our limits on quartic couplings are always applicable, at least numerically.
We classify the states by the conserved quantum numbers $Q$, $Y$, and $T$ and show that, once one restricts oneself to the minimal set of states providing all inequivalent bounds, the quantum number $Q$ is redundant. We also discuss examples where the existence of extra symmetries allows for the inclusion of further quantum numbers in the basis, thus greatly simplifying the matrices.
@@ Appendix A Redundant scattering matrices
In any scattering process, conserved quantum numbers constrain the possible initial and final states. In a globally symmetric $SU(2)\times U(1)$ quantum field theory, initial states with definite electric charge $Q$ and hypercharge $Y$ can only scatter into states with the same $Q$ and $Y$. That is because these quantum numbers are protected by the $SU(2)\times U(1)$ symmetry and are, thus, conserved. For this reason, in building $2\to 2$ scattering matrices, Ref. Bento:2017eti labeled two-particle states by $|Q,Y\rangle$. For models like the one defined in Section II, we list in Table. 9 all possible two-particle states.
@@
$\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\delta_{ca,db},$ (146)
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\phi^{0}_{b}\to\phi^{0}_{c}\phi^{0}_{d}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (147)
@@
Therefore, perturbative unitarity bounds obtained from scatterings involving $|1,\frac{3}{2}\rangle$, $|0,1\rangle$ , and $|0,\frac{1}{2}\rangle$ are redundant because they are identical to those derived from $|2,\frac{3}{2}\rangle$, $|2,1\rangle$ , and $|1,\frac{1}{2}\rangle$ , respectivelly.
In any scattering involving $SU(2)$ doublets, total isospin $T$ must also be conserved, For this reason, two-particle states should be further labeled by $|Q,Y,T\rangle$, with $T=0,1$ in this model. Using Clebsch-Gordan coefficients, we can split the two-particle state $\phi^{+}_{i}\phi^{0}_{j}$ in $|1,1\rangle$ into
@@
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{(a}\,\phi^{0}_{b)}\to\phi^{+}_{(c}\,\phi^{0}_{d)}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (155)
@@
We therefore conclude that it is sufficient to apply the partial wave unitarity bounds to the scattering matrices built out of the states listed in Table 1, i.e. the states
@@ Appendix B The Standard Model example
Table 10: Basis of two-particle states labelled by $|Q,Y\rangle$.
@@
Table 11: Basis of two-particle states labelled by $|Q,Y,T\rangle$.
@@ Appendix C BounDS
We have developed BounDS, a Mathematica notebook, that automates the process of writing the quartic potential and deriving partial wave unitarity bounds for models within the class of models defined in Section II. Notice that ScannerS Coimbra:2013qq ; Muhlleitner:2020wwk has a tool to calculate the scattering matrices, but the user must introduce the correct potential. Here, the user just introduces the number of fields and their symmetries; our program calculates the correct quartic potential and then the relevant scattering matrices. As stated, the notebook can be downloaded from
@@
The list Assume contains information about group theoretical parameters, such as the space they are defined in and relationships between them. For these parameters, the user must use Latin characters. Furthermore, the function Conjugate[] may be used for symmetries that require conjugation of a field. BounDS has been successfully tested with a wide range of symmetry groups, including both discrete and continuous, abelian and non-abelian cases.
@@
Finally, step 3 is dedicated to visualizing and analyzing the output. By calling Potential4, the user can output the quartic part of the most general scalar potential allowed by the symmetries of the model. Anywhere in the code, the Mathematica function TeXForm can be called to get the source code:
@@ Appendix D $\mathbb{Z}_{2}$-Symmetric 2HDM
In a process involving complex scalars that are singlets under the SM gauge group, there are CP-even and CP-odd components that scatter independently. Furthermore, if there are additional flavour symmetries, each two-particle state must also be labelled by its corresponding charge $S_{\alpha}$, following the idea of Ginzburg:2005dt . Therefore, we can label all states by,
@@
Table 12: Basis of two-particle states labelled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$.
@@
We find that, by labelling states with flavour symmetries in addition to charge, hypercharge, and total isospin, the scattering matrices simplify further and reduce in dimensionality compared to those in Section V.3.1. This reduction greatly facilitates the subsequent calculations, as shown in this appendix.

### v3

@@ V.9.1 Scattering Matrices
$A$ $=$ $\begin{bmatrix}2(\lambda_{11,33}+\lambda_{13,31})&2\sqrt{2}\lambda_{12,32}\\[2.84526pt] 2\sqrt{2}\lambda_{12,32}^{*}&2\lambda_{22,22}\end{bmatrix}$, (131)
$B$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{12,13}\\[2.84526pt] 2\sqrt{2}\lambda_{12,13}^{*}&2(\lambda_{22,33}+\lambda_{23,32})\end{bmatrix}$, (132)
$C$ $=$ $\begin{bmatrix}2\lambda_{33,33}&2\sqrt{2}\lambda_{13,23}^{*}\\[2.84526pt] 2\sqrt{2}\lambda_{13,23}&2(\lambda_{11,22}+\lambda_{12,21})\end{bmatrix}$, (133)
@@
$D$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{13,23}^{*}&2\lambda_{12,32}\\ 2\lambda_{13,23}&2\lambda_{11,33}&2\lambda_{12,13}\\ 2\lambda_{12,32}^{*}&2\lambda_{12,13}^{*}&2\lambda_{11,22}\end{bmatrix}$, (135)
@@
$F$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{12,32}^{*}&2\lambda_{13,23}\\ 2\lambda_{12,32}&2\lambda_{11,22}&2\lambda_{12,13}\\ 2\lambda_{13,23}^{*}&2\lambda_{12,13}^{*}&2\lambda_{11,33}\end{bmatrix}$, (137)
@@
$G$ $=$ $\begin{bmatrix}2(\lambda_{22,33}+2\lambda_{23,32})&6\lambda^{*}_{13,23}&6\lambda_{12,32}\\ 6\lambda_{13,23}&2(\lambda_{11,33}+2\lambda_{13,31})&6\lambda_{12,13}\\ 6\lambda^{*}_{12,32}&6\lambda^{*}_{12,13}&2(\lambda_{11,22}+2\lambda_{12,21})\end{bmatrix}$, (139)
$H$ $=$ $\begin{bmatrix}6\lambda_{11,11}&2(2\lambda_{11,22}+\lambda_{12,21})&2(2\lambda_{11,33}+\lambda_{13,31})\\ 2(2\lambda_{11,22}+\lambda_{12,21})&6\lambda_{22,22}&2(2\lambda_{22,33}+\lambda_{23,32})\\ 2(2\lambda_{11,33}+\lambda_{13,31})&2(2\lambda_{22,33}+\lambda_{23,32})&6\lambda_{33,33}\end{bmatrix}$, (140)
$I$ $=$ $\begin{bmatrix}2(\lambda_{22,33}+2\lambda_{23,32})&6\lambda^{*}_{12,32}&6\lambda_{13,23}\\ 6\lambda_{12,32}&2(\lambda_{11,22}+2\lambda_{12,21})&6\lambda_{12,13}\\ 6\lambda^{*}_{13,23}&6\lambda^{*}_{12,13}&2(\lambda_{11,33}+2\lambda_{13,31})\end{bmatrix}$, (141)
@@
V.9.2 Perturbative Unitarity Bounds
We show the results of partial-wave unitarity bounds to the eigenvalues that can be computed analytically,
@@ V.9.2 Perturbative Unitarity Bounds
$\left|\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\pm\sqrt{8|\lambda_{12,13}|^{2}+\left(\lambda_{22,33}+\lambda_{23,32}-\lambda_{11,11}\right){}^{2}}\right|$ $\leq 8\pi,$ (146)
$\left|\lambda_{22,22}+\lambda_{11,33}+\lambda_{13,31}\pm\sqrt{8|\lambda_{12,32}|^{2}+\left(\lambda_{11,33}+\lambda_{13,31}-\lambda_{22,22}\right){}^{2}}\right|$ $\leq 8\pi,$ (147)
$\left|\lambda_{33,33}+\lambda_{11,22}+\lambda_{12,21}\pm\sqrt{8|\lambda_{13,23}|^{2}+\left(\lambda_{11,22}+\lambda_{12,21}-\lambda_{33,33}\right){}^{2}}\right|$ $\leq 8\pi.$ (148)
Table 8: Comparison of coupling notation.
Term Our Notation Notation of Bento et al. (2017)
@@
We compared our results with those presented in Bento et al. (2017), and found them to be consistent. Table 8 provides a summary of the correspondence between the quartic couplings used in our study and those in Bento et al. (2017).
@@ VI Overview
Almost all models addressing the outstanding issues in the SM include extra $SU(2)$ singlet and/or doublet scalars. In particular, many models addressing the dark matter problem include extra neutral singlet scalars. Such models must be subject to theoretical constraints, even before a simulation starts. Those constraints include boundedness from below, nonexistence of lower lying alternative vacua, and the perturbative unitarity bounds on $2\rightarrow 2$ scattering. In this paper, we address the perturbative partial-wave unitarity for the tree-level scattering matrix in models with any number of scalar doublets, neutral singlets, and/or charged singlets. Enforcing the correct high-energy behavior provides bounds on the quartic couplings, freeing us from defining the exact nature of the quadratic and cubic couplings.
In contrast, if one wishes to turn such bounds into restrictions on masses, mixing angles and other directly observable quantities, then one must define the full theory. That is, one must define the quadratic and cubic couplings, a specific vacuum, and the mass matrices must be duly diagonalized. After this, one would strive to invert the relations, turning the unitarity bounds on the quartic couplings into restrictions on combinations of masses and mixing angles. This is possible in simple models, but increasingly more difficult as the number of fields increases. In contrast, although the matrices get larger, our limits on quartic couplings are always applicable, at least numerically.
We classify the states by the conserved quantum numbers $Q$, $Y$, and $T$, and show that, once one restricts oneself to the minimal set of states providing all inequivalent bounds, the quantum number $Q$ is redundant. We also discuss examples where the existence of extra symmetries allows for the inclusion of further quantum numbers in the basis, thus greatly simplifying the scattering matrices.
@@ Appendix A Redundant scattering matrices
In any scattering process, conserved quantum numbers constrain the possible initial and final states. In a globally symmetric $SU(2)\times U(1)$ quantum field theory, initial states with definite electric charge $Q$ and hypercharge $Y$ can only scatter into states with the same $Q$ and $Y$. That is because these quantum numbers are protected by the $SU(2)\times U(1)$ symmetry and are, thus, conserved. For this reason, in building $2\to 2$ scattering matrices, Ref. Bento et al. (2017) labeled two-particle states by $|Q,Y\rangle$. For models like the one defined in Section II, we list in Table. 9 all possible two-particle states.
@@
$\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\varphi^{+}_{b}\to\phi^{0}_{c}\varphi^{+}_{d}\right]=\delta_{ca,db},$ (149)
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\phi^{0}_{b}\to\phi^{0}_{c}\phi^{0}_{d}\right]=2\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (150)
@@
Therefore, perturbative unitarity bounds obtained from scatterings involving $|1,\frac{3}{2}\rangle$, $|0,1\rangle$ , and $|0,\frac{1}{2}\rangle$ are redundant because they are identical to those derived from $|2,\frac{3}{2}\rangle$, $|2,1\rangle$ , and $|1,\frac{1}{2}\rangle$ , respectively.
In any scattering involving $SU(2)$ doublets, total isospin $T$ must also be conserved, For this reason, two-particle states should be further labeled by $|Q,Y,T\rangle$, with $T=0,1$ in this class of models. Using Clebsch-Gordan coefficients, we can split the two-particle state $\phi^{+}_{i}\phi^{0}_{j}$ in $|1,1\rangle$ into
@@
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{(a}\,\phi^{0}_{b)}\to\phi^{+}_{(c}\,\phi^{0}_{d)}\right]=2\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (158)
@@
We therefore conclude that it is sufficient to apply the partial-wave unitarity bounds to the scattering matrices built out of the states listed in Table 1, i.e. the states
@@ Appendix B The Standard Model example
Table 10: Basis of two-particle states labeled by $|Q,Y\rangle$.
@@
Table 11: Basis of two-particle states labeled by $|Q,Y,T\rangle$.
@@ Appendix C BounDS
We have developed BounDS, a Mathematica notebook, that automates the process of writing the quartic potential and deriving partial-wave unitarity bounds for models within the class of models defined in Section II. Notice that ScannerS Coimbra et al. (2013); Mühlleitner et al. (2022) has a tool to calculate the scattering matrices, but the user must introduce the correct potential. Here, the user just introduces the number of fields and their symmetries; our program calculates the correct quartic potential and then the relevant scattering matrices. As stated, the notebook can be downloaded from
@@
The list Assume contains information about group theoretical parameters, such as the space they are defined in and relationships between them. Furthermore, the function Conjugate[] may be used for symmetries that require conjugation of a field. BounDS has been successfully tested with a wide range of symmetry groups, including both discrete and continuous, Abelian and non-Abelian cases. The list Sym[] is read sequentially, and some symmetries may take longer to resolve than others. For improved time performance, we recommend that the user declare Abelian symmetries before imposing possible non-Abelian symmetries.
@@
Finally, step 3 is dedicated to visualizing and analyzing the output. By calling Potential4, the user can output the quartic part of the most general scalar potential allowed by the symmetries of the model. Anywhere in the code, the native Mathematica function TeXForm can be called to get the source code:
@@ Appendix D $\mathbb{Z}_{2}$-Symmetric 2HDM
In a process involving complex scalars that are singlets under the SM gauge group, there are CP-even and CP-odd components that scatter independently. Furthermore, if there are additional flavour symmetries, each two-particle state must also be labeled by its corresponding charge $S_{\alpha}$, following the idea of Ginzburg and Ivanov (2005). Therefore, we can label all states by,
@@
Table 12: Basis of two-particle states labeled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$.
@@
We find that, by labeling states with flavour symmetries in addition to charge, hypercharge, and total isospin, the scattering matrices simplify further and reduce in dimensionality compared to those in Section V.3.1. This reduction greatly facilitates the subsequent calculations, as shown in this appendix.